# Intuitionistic NF set theory

### v5

@@
Abstract.
We develop NF set theory using intuitionistic logic; we call this theory INF. We develop the theories of finite sets and their power sets, finite cardinals and their ordering, cardinal exponentiation, addition, and multiplication. We follow Rosser and Specker with appropriate constructive modifications, especially replacing “arbitrary subset” by “separable subset” in the definitions of exponentiation and order. It is not known whether INF proves that the set of finite cardinals is infinite, so the whole development must allow for the possibility that there is a maximum integer; arithmetical computations might “overflow” as in a computer or odometer, and theorems about them must be carefully stated to allow for this possibility. The work presented here is intended as a substrate for further investigations of INF.
1. Introduction
Quine’s NF set theory is a first-order theory whose language contains only the binary predicate symbol $\in$, and whose axioms are two in number: extensionality and stratified comprehension. The definition of these axioms will be reviewed below; full details can be found in [10]. Intuitionistic NF, or INF, is the theory with the same language and axioms as NF, but with intuitionistic logic instead of classical.
The “axiom” of infinity is a theorem of NF, proved by Rosser [8, 10] and Specker [11]. These proofs use classical logic in an apparently essential way. It is still an open question whether INF proves the existence of an infinite set. The Stanford Encyclopedia of Philosophy’s article on NF says [3]
The only known proof (Specker’s) of the axiom of infinity in NF has too little constructive content to allow a demonstration that INF admits an implementation of Heyting arithmetic.
This paper is devoted to supplying infrastructure, on the basis of which one can investigate INF. Here is a list of open problems about INF: [^1] The terms used in these questions can be looked up in the index of [1].
@@ 1. Introduction
  • Can one point to any specific instance of the law of the excluded middle that is not provable in INF (even assuming NF is consistent)?
@@
  • Is INF consistent? Is there any double-negation interpretation from NF to INF?
@@
  • Is Church’s thesis consistent with INF? Markov’s principle?
  • Is INF closed under Church’s rule?
@@
Regarding the problem whether ${\mathbb{F}}$ is finite: For all we know, there might be a largest finite cardinal $m$, which would contain a finite set $U$ that is “unenlargeable”, in the sense that we cannot find any $x$ that is not a member of $U$. Classically, that would imply $U={\mathbb{V}}$, which is a contradiction, since ${\mathbb{V}}$ is not finite. But intuitionistically, it is an open question.
Thanks to Thomas Forster for asking me (once a year for twenty years) about the strength of INF. Thanks to Randall Holmes, Albert Visser, and Thomas Forster for many emails on this subject. Thanks to the creators of the proof assistant Lean [2], which has enabled me to state with high confidence that there are no errors in this paper.222 Each proof in this paper has been formalized and checked for validity using the proof assistant Lean. However, the proofs presented in this paper are human-readable and as complete as a detailed human-readable proof needs to be. Thanks to the users of Lean who helped me acquire sufficient expertise in using Lean by answering my questions, especially Mario Carneiro.
There are many lemmas in this paper, and each of those lemmas is provable in INF. Occasionally we state explicitly “INF proves”, but even when we do not, it is the case. An important reference for NF is Rosser’s book [9, 10].333 The two editions are identical except for the Appendices added to the second edition, one of which contains Rosser’s proof of infinity. We follow some of Rosser’s notation, e.g., $\Lambda$ for the empty set, and $Nc(x)$ for the cardinal of $x$, $USC(x)$ for the class of unit subclasses (singletons) of $x$, ${\mathbb{V}}$ for the universe.
@@
But the logical apparatus of Rosser’s system includes a Hilbert-style epsilon-operator, which is not compatible with an intuitionistic version, and also, we do not wish to assume the axiom of infinity. Since all of Rosser’s results are obtained using classical logic, we must begin afresh, proving exactly what we need in a first-order intuitionistic system.
@@
It should be noted that the consistency of classical NF has been proved [6], and the proof has been checked in Lean. This result implies, of course, that the subtheory INF considered here is also consistent; but is otherwise not directly relevant.
2. Axioms of NF, ordered pairs, and functions
@@
NF has exactly two axioms: extensionality and stratified comprehension. The axiom of extensionality says that two sets with the same elements are equal. The axiom of stratified comprehension says that $\{x:\phi(x)\}$ exists, if $\phi$ is a stratified formula. A formula is stratified, or stratifiable, if each of its variables (both bound and free) can be assigned a non-negative integer (“index” or “type”) such that (i) in every subformula $x\in y$, $y$ gets an index one greater than $x$ gets, and (ii) every occurrence of each variable gets the same index.
@@
Proof. Straightforward application of the definition and extensionality. We omit the approximately 70-step proof.
@@
Then a function of two variables is definable in INF if its graph forms a set of ordered triples $\langle x,y,f(x,y)\rangle$.
We can conservatively add function symbols for binary union $x\cup y$ and intersection $x\cap y$, union, intersection, set difference $x-y$, and generally we can add a function symbol $c_{\phi}$ for any stratified formula $\phi$, so that $x\in c_{\phi}(y){\ \leftrightarrow\ }\phi(x,y)$. For a detailed discussion of the logical underpinnings of this step, see [5]. Function symbols for $\{x\}$, $\{x,y\}$, and $\langle x,y\rangle$ are also special cases of the $c_{\phi}$; we can add these function symbols even though the “functions” they denote are not functions in the sense that their graphs are definable in INF. Thus for example we have
@@
Proof. This is the defining axiom for the function symbol $\{x\}$, which is really just $\{u:u=x\}$; that is, the function symbol is $c_{\phi}$ where $\phi(u,x)$ is $u=x$.
@@
Proof. This can be taken as the defining axiom for $a-b$; or it may be derived in a finite axiomatization from other axioms.
@@
Proof. Right to left is just equality substitution. Ad left to right: Suppose $\{x\}=\{y\}$. Then
@@
In practice we need to use stratified comprehension in the presence of function symbols and parameters; the notion of stratification has to be extended to cover these situations. We define the notion of a formula $\phi$ being “stratified with respect to $x$”. The variables of $\phi(x)$ are of three kinds: $x$ (the “eigenvariable”), variables other than $x$ that occur only on the right of $\in$ (“parameters”), and all other variables. An assignment of natural numbers (indices) to the variables that are not parameters is said to stratify $\phi$ with respect to $x$ if for each atomic formula $z\in y$, $y$ is assigned an index one larger than the index assigned to $z$. Note that the assignment is to variables, rather than occurrences of variable, so every occurrence of $z$ gets the same index. Note also that parameters need not be assigned an index.
@@
It is “well-known” that stratified comprehension, so defined, is conservative over NF, but it does not seem to proved in the standard references on NF; and besides, we need that result for INF as well. The algorithm in [5] meets the need: it will unwind the function symbols in favor of their definitions, preserving stratification. The confused reader is advised to work this out on paper for the example of the binary function symbol $\langle x,y\rangle$.
In our work, we repeatedly assert that certain formulas are stratifiable, and then we apply comprehension, either directly or indirectly by using mathematical induction or induction on finite sets. The question then arises of ensuring that only correctly stratified instances of comprehension are used. One approach is to use a finite axiomatization of INF. (It is easy to write one down following well-known examples for classical NF.) But that just pushes the problem back to verifying the correctness of that axiomatization; moreover it is technically difficult to reduce given particular instances of comprehension to a finite axiomatization. Instead, we just made a list of each instance of comprehension that we needed. There are 154 instances of comprehension in that list (which includes more than just the instances used in this paper). The Lean proof assistant does not check that those instances are stratified. If one is not satisfied with a manual check of those 154 formulas, then one has to write a computer program to check that they are stratified. We did write one and those 154 formulas passed; since this paper is being presented as human-readable, we rely here on the human reader to check each stratification as it is presented; we shall not go into the technicalities of computer-checking stratification.
@@
$f:X\to Y$ (“$f$ maps $X$ to $Y$”) means for every $x\in X$ there exists a unique $y\in Y$ such that $\langle x,y\rangle\in f$. “$f$ is a function” means
@@
When $f$ is a function, one writes $f(x)$ for that unique $y$. It is time to justify that practice in the context of INF. [^5] Rosser’s version of classical NF has Hilbert-style choice operator, which gives us “some $y$ such that $\langle x,y\rangle\in f$.” But INF does not and cannot have such an operator, so a different formal treatment is needed. Here is how to do that. We introduce a function symbol $Ap$ (with the idea that we will abbreviate $Ap(f,x)$ to $f(x)$ informally).
@@
Proof. Suppose $f$ is a function and $\langle x,y\rangle\in f$. We must prove $y=Ap(f,x)$. By extensionality it suffices to show that for all $t$,
(1) $t\in y{\ \leftrightarrow\ }t\in Ap(f,x)$
@@
Right to left: Suppose $t\in Ap(f,x)$. Then by the definition of $Ap$, for some $z$ we have $\langle x,z\rangle\in f$ and $t\in z$. Since $f$ is a function, $y=z$. Then $t\in y$. That completes the right-to-left direction. That completes the proof of the lemma.
@@
The definition of $f^{-1}$ can be given by a stratified formula, so it is legal in INF.
@@
Proof. Let $f:X\to Y$ be one-to-one and onto. Since $f$ is one-to-one, for each $y\in Y$ there is a unique $x$ such that $\langle x,y\rangle\in f$. Then by definition of function, $f^{-1}:y\to x$. Since $f:X\to Y$, for each $x\in X$ there is a unique $y\in Y$ such that $\langle x,y\rangle\in f$.
@@
I say $f^{-1}$ maps $Y$ onto $X$. Let $x\in X$. Let $y=f(x)$. Then $\langle x,y\rangle\in f$. Then $\langle y,x\rangle\in f^{-1}$. Therefore $f^{-1}$ is onto, as claimed. That completes the proof of the lemma.
@@
The relation “$x$ is similar to $y$” is defined by
[…]
@@
The defining formula is stratified giving $x$ and $y$ the same type, so the relation is definable in INF.
@@
Proof. Ad reflexivity: $x\sim x$ because the identity map from $x$ to $x$ is one-to-one and onto.
@@
Ad transitivity: Let $x\sim y$ and $y\sim z$. Then there exist $f$ and $g$ such that $f:x\to y$ is one-to-one and onto, and $g:y\to z$ is one-to-one and onto. Then $f\circ g:x\to z$ is one-to-one and onto. Therefore $x\sim z$. That completes the proof of transitivity. That completes the proof of the lemma.
@@
$x\sim\Lambda{\ \leftrightarrow\ }x=\Lambda.$
Proof. Left to right: suppose $x\sim\Lambda$. Let $f:x\to\Lambda$ be a similarity. Suppose $u\in x$. Then for some $v$, $\langle u,v\rangle\in f$ and $v\in\Lambda$. But $v\not\in\Lambda$. Hence $u\not\in x$. Since $u$ was arbitrary, $x=\Lambda$, as desired.
Right to left: Suppose $x=\Lambda$. We have to show $\Lambda\sim\Lambda$. But $\Lambda:\Lambda\to\Lambda$ is a similarity. That completes the proof of the lemma.
@@
Proof. By the definition of $\subseteq$ and the axiom of extensionality.
3. Finite sets

### v6

# Finite sets, mappings, cardinals, and arithmetic in intuitionistic NF

@@
Abstract
NF set theory using intuitionistic logic is called $\mathit{i}\mkern 0.4mu\mathrm{NF}$. We develop the theories of finite sets and their power sets and mappings, finite cardinals and their ordering, cardinal exponentiation, addition, and multiplication. We follow Rosser and Specker with appropriate constructive modifications, especially replacing “arbitrary subset” by “separable subset” in the definitions of exponentiation and order. It is not known whether $\mathit{i}\mkern 0.4mu\mathrm{NF}$ proves that the set of finite cardinals is infinite, so the whole development must allow for the possibility that there is a maximum integer; arithmetical computations might “overflow” as in a computer or odometer, and theorems about them must be carefully stated to allow for this possibility. The work presented here is intended as a substrate for further investigations of $\mathit{i}\mkern 0.4mu\mathrm{NF}$, including the development of Bishop-style constructive mathematics in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
1 Introduction
[…]
The “axiom” of infinity is a theorem of NF, proved by Rosser  [14, 16] and Specker [17]. These proofs use classical logic in an apparently essential way. It is still an open question whether $\mathit{i}\mkern 0.4mu\mathrm{NF}$ proves the existence of an infinite set. The Stanford Encyclopedia of Philosophy article on NF says [6]
The only known proof (Specker’s) of the axiom of infinity in NF has too little constructive content to allow a demonstration that $\mathit{i}\mkern 0.4mu\mathrm{NF}$ admits an implementation of Heyting arithmetic.
In attempting to determine whether the quoted statement is true, I found that I first needed to develop enough basic mathematics in $\mathit{i}\mkern 0.4mu\mathrm{NF}$ to tackle Specker’s proof. That mathematical infrastructure is presented in this paper. [^3] Eventually I came to the conclusion that the quoted statement is true; that is beyond the scope of this paper, but see a short discussion near the end. The purpose of this development is to provide a basis on which one can:
  • investigate $\mathit{i}\mkern 0.4mu\mathrm{NF}$ further;
  • develop Bishop-style constructive mathematics in $\mathit{i}\mkern 0.4mu\mathrm{NF}$
    (after proving or assuming infinity).
The following questions about $\mathit{i}\mkern 0.4mu\mathrm{NF}$ remain open: [^4] The terms used in these questions can be looked up in the index of [1].
@@ 1 Introduction
  • Can one point to any specific instance of the law of the excluded middle that is not provable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$ (even assuming NF is consistent)?
  • Is $\mathit{i}\mkern 0.4mu\mathrm{NF}$ consistent?
@@
  • Is there any double-negation interpretation from NF to $\mathit{i}\mkern 0.4mu\mathrm{NF}$?
@@
  • Is Church’s thesis consistent with $\mathit{i}\mkern 0.4mu\mathrm{NF}$? Markov’s principle?
  • Is $\mathit{i}\mkern 0.4mu\mathrm{NF}$ closed under Church’s rule?
@@
Regarding whether $\mathbb{F}$} is finite: For all we know, there might be a largest finite cardinal $\bf m$}, which would contain a finite set $U$ that is “unenlargeable”, in the sense that we cannot find any $x$ that is not a member of $U$. Classically, that would imply $U={\mathbb{V}}$, which is a contradiction, since $\mathbb{V}$} is not finite. But intuitionistically, it is an open question.
@@
Each of the lemmas and theorems in this paper is provable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. An important reference for NF is Rosser’s book [15, 16]. [^5] The two editions are identical except for the Appendices added to the second edition, one of which contains Rosser’s proof of infinity. But the logical apparatus of Rosser’s system includes a Hilbert-style epsilon-operator, which is not compatible with an intuitionistic version, and also, we do not wish to assume the axiom of infinity. Since all of Rosser’s results are obtained using classical logic, we cannot rely on Rosser.
@@
It should be noted that the consistency of classical NF has been proved [12], and the proof has been checked in Lean. This result implies, of course, that the subtheory $\mathit{i}\mkern 0.4mu\mathrm{NF}$ considered here is also consistent; but it is otherwise not directly relevant.
Notational issues. There is no traditional, universally accepted notation for some of the notions central to NF. Rosser [16] and Specker [17] are two of the original sources. Both of these were written prior to the advent of and , and for the most part were limited to characters found on a typewriter keyboard. Forster [9] used different notation, making use of . Now, however, keyboard characters are back in style, because they are easier to use in computer proof-checking. Lean, for example, goes to great lengths to support complicated typography—but one cannot search for those symbols, which is quite annoying.
I therefore used Specker’s notation when using Lean and in pre-publication versions of this paper; but the referee asked me to change it, so I did. The following table compares the notational styles of Specker and this paper. It may prove useful if anyone wants to compare this paper to Specker or Rosser, or to the Lean proofs, or to readers who know one style or the other already.
Table 1: Notation
Specker This paper
$USC(x)$ $\mathcal{P}_{1}(x)$
$SSC(x)$ $\mathcal{P}_{s}(x)$
$SC(x)$ $\mathcal{P}(x)$
$Nc(x)$ $|x|$
$\Lambda$ $\emptyset$
Use of computer proof-checking. All the proofs in this paper have been checked in the proof assistant Lean. Could there still be errors? The possible sources of error are
  • Use of an unstratified definition
  • Lean proof and paper proof might not exactly correspond
  • Lean might have smuggled in classical logic, i.e., used it without telling me.
  • Perhaps the order of theorems in the paper is not strictly the logical order.
Regarding the smuggling: Lean’s underlying theorem is intuitionistic, but the library is classical, and even though I didn’t use the library, and even though Lean experts helped me, the possibility theoretically exists. Regarding the correspondence: if there are such problems, they are just typos. Regarding stratification: I allowed the full comprehension axiom, but used only stratified instances. I used a computer script ex post facto to check stratification. [^6] Originally I intended to use a finite axiomatization. But it is often quite complicated to derive simple definitions from a finite axiomatization; and then one still has to worry if the finite axiomatization is really correct. Regarding the order of theorems: At least no lemma or theorem is cited before it is proved. Of course in Lean, the logical order is enforced, but that is often not the best order for presentation.
The reader who is worried about errors in Lean has the option to forget it was ever mentioned, and just read the proofs, which are here presented in complete human-readable detail.
Acknowledgements. Thanks to Thomas Forster for asking me (once a year for twenty years) about the strength of $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Thanks to Randall Holmes, Albert Visser, and Thomas Forster for many emails on this subject. Thanks to the creators of the proof assistant Lean [2], which has enabled me to state with high confidence that there are no errors in this paper. Thanks to the users of Lean who helped me acquire sufficient expertise in using Lean by answering my questions, especially Mario Carneiro.
2 Axioms of NF, ordered pairs, and functions
@@
NF has exactly two axioms: extensionality and stratified comprehension. The axiom of extensionality says that two sets with the same elements are equal. The axiom schema of stratified comprehension says that $\{x:\phi(x)\}$ exists, if $\phi$ is a stratified formula. A formula is stratified, or stratifiable, if each of its variables (both bound and free) can be assigned a non-negative integer (“index” or “type”) such that (i) in every subformula $x\in y$, $y$ gets an index one greater than $x$ gets, and (ii) every occurrence of each variable gets the same index.
@@
Proof.
Straightforward application of the definition and extensionality. We omit the approximately 70-step proof. ∎
@@
Then a function of two variables is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$ if its graph forms a set of ordered triples $\langle x,y,f(x,y)\rangle$.
We can conservatively add function symbols for binary union $x\cup y$ and intersection $x\cap y$, union, intersection, set difference xyx-y, and generally we can add a function symbol $c_{\phi}$ for any stratified formula $\phi$, so that $x\in c_{\phi}(y){\ \leftrightarrow\ }\phi(x,y)$. For a detailed discussion of the logical underpinnings of this step, see [10]. Function symbols for $\{x\}$, $\{x,y\}$, and $\langle x,y\rangle$ are also special cases of the $c_{\phi}$; we can add these function symbols even though the “functions” they denote are not functions in the sense that their graphs are definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Thus for example we have
@@
Proof.
This is the defining axiom for the function symbol $\{x\}$, which is really just $\{u:u=x\}$; that is, the function symbol is $c_{\phi}$ where $\phi(u,x)$ is u=xu=x.∎
@@
Proof.
This can be taken as the defining axiom for aba-b; or it may be derived in a finite axiomatization from other axioms. ∎
@@
Proof.
Right to left is just equality substitution. Ad left to right: Suppose $\{x\}=\{y\}$. Then
@@
In practice we need to use stratified comprehension in the presence of function symbols and parameters; the notion of stratification has to be extended to cover these situations. We define the notion of a formula $\phi$ being “stratified with respect to xx”. The variables of $\phi(x)$ are of three kinds: $x$ (the “eigenvariable”), variables other than $x$ that occur only on the right of $\in$ (“parameters”), and all other variables. An assignment of natural numbers (indices) to the variables that are not parameters is said to stratify $\phi$ with respect to $x$ if for each atomic formula $z\in y$, $y$ is assigned an index one larger than the index assigned to $z$. Note that the assignment is to variables, rather than occurrences of variable, so every occurrence of $z$ gets the same index. Note also that parameters need not be assigned an index.
@@
It is “well-known” that stratified comprehension, so defined, is conservative over NF, but it does not seem to proved in the standard references on NF; and besides, we need that result for $\mathit{i}\mkern 0.4mu\mathrm{NF}$ as well. The algorithm in  [10] meets the need: it will unwind the function symbols in favor of their definitions, preserving stratification. The confused reader is advised to work this out on paper for the example of the binary function symbol $\langle x,y\rangle$.
In our work, we repeatedly assert that certain formulas are stratifiable, and then we apply comprehension, either directly or indirectly by using mathematical induction or induction on finite sets. The question then arises of ensuring that only correctly stratified instances of comprehension are used. One approach is to use a finite axiomatization of $\mathit{i}\mkern 0.4mu\mathrm{NF}$. (It is easy to write one down following well-known examples for classical NF.) But that just pushes the problem back to verifying the correctness of that axiomatization; moreover it is technically difficult to reduce given particular instances of comprehension to a finite axiomatization. Instead, we just made a list of each instance of comprehension that we needed. There were at some point 154 instances of comprehension in that list (which includes more than just the instances used in this paper). The Lean proof assistant does not check that those instances are stratified. If one is not satisfied with a manual check of those 154 formulas, then one has to write a computer program to check that they are stratified. We did write one and those 154 formulas passed; since this paper is being presented as human-readable, we rely here on the human reader to check each stratification as it is presented; we shall not go into the technicalities of computer-checking stratification.
@@
$f:X\to Y$ (“$f$ maps $X$ to YY”) means for every $x\in X$ there exists a unique $y\in Y$ such that $\langle x,y\rangle\in f$. “$f$ is a function” means
@@
When $f$ is a function, one writes $f(x)$ for that unique $y$. It is time to justify that practice in the context of $\mathit{i}\mkern 0.4mu\mathrm{NF}$. [^8] Rosser’s version of classical NF has Hilbert-style choice operator, which gives us “some $y$ such that $\langle x,y\rangle\in f$.” But $\mathit{i}\mkern 0.4mu\mathrm{NF}$ does not and cannot have such an operator, so a different formal treatment is needed. Here is how to do that. We introduce a function symbol $Ap$ (with the idea that we will abbreviate $Ap(f,x)$ to $f(x)$ informally).
@@
Proof.
Suppose $f$ is a function and $\langle x,y\rangle\in f$. We must prove $y=Ap(f,x)$. By extensionality it suffices to show that for all $t$,
$t\in y{\ \leftrightarrow\ }t\in Ap(f,x)$ (1)
@@
Right to left: Suppose $t\in Ap(f,x)$. Then by the definition of $Ap$, for some $z$ we have $\langle x,z\rangle\in f$ and $t\in z$. Since $f$ is a function, $y=z$. Then $t\in y$. That completes the right-to-left direction. ∎
@@
The definition of $f^{-1}$ can be given by a stratified formula, so it is legal in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
@@
Proof.
Let $f:X\to Y$ be one-to-one and onto. Since $f$ is one-to-one, for each $y\in Y$ there is a unique $x$ such that $\langle x,y\rangle\in f$. Then by definition of function, $f^{-1}:y\to x$. Since $f:X\to Y$, for each $x\in X$ there is a unique $y\in Y$ such that $\langle x,y\rangle\in f$.
@@
I say $f^{-1}$ maps $Y$ onto $X$. Let $x\in X$. Let $y=f(x)$. Then $\langle x,y\rangle\in f$. Then $\langle y,x\rangle\in f^{-1}$. Therefore $f^{-1}$ is onto, as claimed. ∎
@@
The relation “$x$ is similar to yy” is defined by
$x\sim y{\ \leftrightarrow\ }\exists f\,(f:x\to y\ \land\mbox{\, $f$ is one-to-one and onto}).$
@@
The defining formula is stratified giving $x$ and $y$ the same type, so the relation is definable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
@@
Proof.
Ad reflexivity: $x\sim x$ because the identity map from $x$ to $x$ is one-to-one and onto.
@@
Ad transitivity: Let $x\sim y$ and $y\sim z$. Then there exist $f$ and $g$ such that $f:x\to y$ is one-to-one and onto, and $g:y\to z$ is one-to-one and onto. Then $f\circ g:x\to z$ is one-to-one and onto. Therefore $x\sim z$. That completes the proof of transitivity. ∎
@@
$x\sim\emptyset{\ \leftrightarrow\ }x=\emptyset.$
Proof.
Left to right: suppose $x\sim\emptyset$. Let $f:x\to\emptyset$ be a similarity. Suppose $u\in x$. Then for some $v$, $\langle u,v\rangle\in f$ and $v\in\emptyset$. But $v\not\in\emptyset$. Hence $u\not\in x$. Since $u$ was arbitrary, $x=\emptyset$, as desired.
Right to left: Suppose $x=\emptyset$. We have to show $\emptyset\sim\emptyset$. But $\emptyset:\emptyset\to\emptyset$ is a similarity. ∎
@@
Proof.
By the definition of $\subseteq$ and the axiom of extensionality. ∎
3 Finite sets